If the index of refraction of glass for red light is 1.45, what is the speed of a photon in the glass?

a. 2.07 x 108 m/s

b. 4.35 x 108 m/s

c. 300,000 km/s

d. 3 x 108 m/s
Answer #1

speed of photon in a medium is given by

V = c/n

Where n is refractive index of medium.

V = (3×108)/1.45 =

= 2.07×108 m/s

option (a)